OpenStudy (anonymous):

Find x. 1. csc(x)=2 2. csc(x)=1

OpenStudy (anonymous):

That is the same as saying sin(x) = 1/2 and sin(x) = 1. Find x for those values.

OpenStudy (anonymous):

Okay thanks. So x =pi/6 and x=pi/4?

OpenStudy (anonymous):

wait i think pi/2 instead of pi/4

OpenStudy (anonymous):

30 degrees and 90 degrees.

OpenStudy (anonymous):

Thank you I got it!